karissafrazier:

Solve the system of equations and choose the correct ordered pair 4x+3y=16 7x+6y= 31 A. (4,0) B. (4,3) C. (1,4) D. (1,3) pls don't give direct answers that means u @Nishio @Faux

Astro:

Do you know what to do first?

karissafrazier:

@astro wrote:
Do you know what to do first?
sadly no, i'm sorry

Astro:

That's okay, Do you know how to solve for x?

YasssPoo:

4x+3y=16 7x+6y=31 y=(16-4x)/3 y=(31-7x)/6 (16-4x)/3 = (31-7x)/6 6(16-4x) = 3(31-7x) 96-24x=93-21x 3=3x x=1 4x+3y=16, x=1 4(1)+3y=16 4+3y=16 3y=12 y=4 x=1, y=4 therefore the answer is B (1,4)

karissafrazier:

...

VIBE:

@VIBE wrote:
Solve for the first variable in one of the equations, then substitute the result into the other equation. Point Form: (1,4) Equation Form: x=1,y=4

lazyking4life:

c 1, 4

Astro:

Guys, Next time can you let me finish explaining before you just give them the answer.

karissafrazier:

@yassspoo wrote:
4x+3y=16 7x+6y=31 y=(16-4x)/3 y=(31-7x)/6 (16-4x)/3 = (31-7x)/6 6(16-4x) = 3(31-7x) 96-24x=93-21x 3=3x x=1 4x+3y=16, x=1 4(1)+3y=16 4+3y=16 3y=12 y=4 x=1, y=4 therefore the answer is B (1,4)
even tho it was not a direct answer right away u shouldn't give direct answers- it'll help me learn

karissafrazier:

well-since there seems to be a disagreement between A and C i'll try to remember how to do this

Astro:

Okay good, Knowing that it is now between A and C what do you think it is?

karissafrazier:

i think maybe it's C

karissafrazier:

but i could be wrong

YasssPoo:

noo

Rival:

possibly

Astro:

That is correct with C

VIBE:

@astro wrote:
Guys, Next time can you let me finish explaining before you just give them the answer.
i didn't just give them the answer.

karissafrazier:

thank you everyone- mostly @Astro

VIBE:

i told them how to do it.

Astro:

@karissafrazier wrote:
thank you everyone- mostly @Astro
Yw 💜

karissafrazier:

@vibe wrote:
@astro wrote:
Guys, Next time can you let me finish explaining before you just give them the answer.
i didn't just give them the answer.
as well as gave me the answer directly instead of letting me work on it and learning

Astro:

@karissafrazier wrote:
@vibe wrote:
@astro wrote:
Guys, Next time can you let me finish explaining before you just give them the answer.
i didn't just give them the answer.
as well as gave me the answer directly instead of letting me work on it and learning
Very True

karissafrazier:

@VIBE you are lucky i don't block you for that

jamescotal:

It's c

Astro:

@jamescotal wrote:
It's c
We have already figured this out

karissafrazier:

how do i get a reply lock on this question?